OpenStudy (anonymous):

An outside circular ring has a circumference of 90 cm. What is the circumference of an inner ring which is 8 cm shorter in radius? Both circles have the same center. 39.8 correct?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

Circumference = 90 Radius = 90/(2*pi) => 90/6.28 = ?

OpenStudy (maheshmeghwal9):

ya it is correct @suzy4321

OpenStudy (maheshmeghwal9):

I know u have calculated outer circle's radius & then subtract 8 & got inner circle's radius. then u found circumference of inner circle by the fomula as:) \[2\pi r\]

OpenStudy (maheshmeghwal9):

& the answer u got \[39.76\approx39.8\] is correct:)

OpenStudy (anonymous):

39.7

OpenStudy (anonymous):

okay thanks everyone

OpenStudy (anonymous):

My drawings with the calculation I showed are more useful I think. Anyway, all is well!!

OpenStudy (anonymous):

GT that is why u got three medals